OpenStudy (anonymous):

Sorry about yesterday, are you still able to help? @DanJS Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

Or anyone @kevindragonlord

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

the solution to the first one is \((-2,3)\) i suppose you could check which other system h as the same solution

OpenStudy (anonymous):

lets try the last one first \[6x - y = -15\\ 3y = 9\]

OpenStudy (anonymous):

\[6\times -2-3=-12-3=-15\checkmark\\ 3\times 3=9\checkmark\]

OpenStudy (anonymous):

looks like a winner

OpenStudy (anonymous):

So you just add (-2,3) into the equation, do the math and see if it comes out correct?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes, if by "add" you mean replace x by -2 and y by 3

OpenStudy (anonymous):

Where did you get (-2,3) though?

OpenStudy (anonymous):

from the original system

OpenStudy (anonymous):

solve \[6x + 2y = -6\\3x - 4y = -18\] by doubling the first one to get \[12x+4y=-12\\ 3x-4y=-18\] add and get \[15x=-30\\ x=-2\]

OpenStudy (anonymous):

Oh okay, thank you! so then 6x - y = -15 3y = 9 is the answer? @satellite73